The length of a rectangular jewelry box in 3 inches more than twice the width. The perimeter is 42 inches. Find the length and width.
I have is written out as:
42= 2(3+2w)+ 2w
42=2*3+2w+2w
42= 6+4w
-6 -6
36= 4w
4 4
9=w
But when I plug the 9 in for w, it does not equal 42.

Answer by rfer(16322)

multiply both values
2(3+2w)=6+4w
so
42=6+4w+2w
42-6=6w
36=6w
36/6=w
w=6
2w+3=15